OpenStudy (anonymous):

which choice is a solution to the system of equations below? 3y=6x+9 x=3y-4 A.(-1,1) B.no solutions C.infinitely many D.(2,1)

OpenStudy (anonymous):

Okay, first you plug in what x is into the first equation

OpenStudy (anonymous):

ugghhh yeaaah what he said.

OpenStudy (anonymous):

3y=6(3y-4)+9 3y=18y-24+9 3y=18y-15 3y-18y=-15 -15y=-15 y=1 x=3y-4 x=3(1)-4=-1 so. x=-1 and y=1

OpenStudy (anonymous):

A.(-1,1)

OpenStudy (anonymous):

i got A i just didnt know if C could be an answer

OpenStudy (anonymous):

c is not the answer , as it as unique solution